# Unveiling the Unexplored Decay Mode of a Light Charged Higgs Boson to an Off-Shell Top Quark and a Bottom Quark

### v1

@@ II Brief review of the light charged Higgs boson in type-I 2HDM
For the theoretical requirements, we enforce conditions ensuring vacuum stability Ivanov:2008cxa ; Barroso:2012mj ; Barroso:2013awa , a bounded-from-below Higgs potential Ivanov:2006yq , tree-level unitarity in scalar-scalar scatterings Branco:2011iw ; Arhrib:2000is , and perturbativity of the Higgs quartic couplings Chang:2015goa . These conditions are evaluated using the public code 2HDMC Eriksson:2009ws . Additionally, we require that the cutoff scale exceed $10{\;{\rm TeV}}$, where the cutoff scale is defined as the energy level at which any of the conditions for tree-level unitarity, perturbativity, or vacuum stability is violated Kim:2023lxc . The evolution of the model parameters via the renormalization group equations is facilitated using the public code 2HDME Oredsson:2018yho ; Oredsson:2018vio .
For the experimental constraints, we incorporated measurements at the 95% confidence level, encompassing inclusive $B$-meson decay into $X_{s}\gamma$ Arbey:2017gmh ; Sanyal:2019xcp ; Misiak:2017bgg and direct search bounds from LEP, Tevatron, and LHC experiments. For the direct search constraints, we have employed the public code HiggsBounds-v5.10.2 Bechtle:2013wla . Notably, our adoption of the Higgs alignment limit ensures that the Higgs precision data at the LHC are inherently satisfied.
@@
In Figure 1, we present the allowed parameter points in the $(M_{H/A},\tan\beta)$ plane for the given $M_{H^{\pm}}=150{\;{\rm GeV}}$. The color code denotes $m_{12}^{2}$. A notable feature is the upper bound on the masses of $H$ and $A$, with $M_{H/A}\lesssim 236{\;{\rm GeV}}$. Although our focus in this paper is on the charged Higgs boson, the relatively low upper bounds on $M_{H/A}$ suggest promising discovery prospects for the heavy neutral Higgs bosons at high-energy colliders. Additionally, we observe lower bounds on $\tan\beta$, specifically $\tan\beta\gtrsim 6$.
@@
Now, let us identify the optimal production mechanism at the LHC for the light charged Higgs boson. LHC searches have primarily focused on its production through top quark decay, $t\rightarrow bH^{\pm}$, followed by $H^{\pm}\rightarrow\tau\nu$ ATLAS:2018gfm ; Sirunyan:2019hkq because the Yukawa couplings of $H^{\pm}$ are proportional to the fermion mass with a common coupling modifier $1/\tan\beta$. Despite comprehensive searches, no new signals have been observed, leading to stringent upper limits on the product of the two branching ratios, $\text{Br}(t\rightarrow bH^{\pm})\text{Br}(H^{\pm}\rightarrow\tau\nu)$. As $\text{Br}(t\rightarrow bH^{\pm})$ depends solely on $\tan\beta$ for the given $M_{H^{\pm}}$, the observed upper bound strictly limits $\text{Br}(t\rightarrow bH^{\pm})$. For $M_{H^{\pm}}\sim 100{\;{\rm GeV}}$, $\text{Br}(t\rightarrow bH^{\pm})$ must be below $\mathcal{O}(10^{-4})$. Consequently, leveraging top quark decay for $H^{\pm}$ production proves ineffective.
@@
Given this constraint, the production channels of $H^{\pm}$ via the decay of heavier Higgs states, such as $H/A\rightarrow H^{\pm}W^{\mp(*)}$, have been extensively investigated Arhrib:2017wmo ; Mondal:2021bxa ; Arhrib:2021xmc ; Cheung:2022ndq ; Bhatia:2022ugu ; Kim:2022hvh ; Kim:2022nmm ; Arhrib:2022inj ; Kim:2023lxc ; Li:2023btx ; Mondal:2023wib . The production of $H$ or $A$ occurs through gluon fusion ($gg\rightarrow H/A$) or associated production ($gg\rightarrow H\rightarrow AZ$, $q\bar{q}\rightarrow Z\rightarrow HZ$). However, the signal rates are sensitive to model parameters, such as $M_{H/A}$ and $\tan\beta$. Furthermore, the allowed parameter space shown in Figure 1 does not permit the on-shell decay $H/A\rightarrow H^{\pm}W^{\mp}$, preventing the effective suppression of backgrounds by exploiting the $W$ boson mass constraint.
@@
A more promising production mechanism is the pair production of charged Higgs bosons via the Drell-Yan process. This channel offers a straightforward and model-independent avenue for $H^{\pm}$ production, as the production cross section is exclusively determined by the mass of the charged Higgs boson. Although this production channel has been studied for the decays $H^{\pm}\rightarrow\tau\nu$ Duarte:2024zeh and $H^{\pm}\rightarrow W^{\pm(*)}\varphi^{0}/A$ Arhrib:2021xmc ; Arhrib:2021yqf ; Shen:2022yuo , it has not been explored for our target decay mode $H^{\pm}\rightarrow t^{*}b$. Therefore, we focus our analysis on the pair production channel for the $H^{\pm}\rightarrow t^{*}b$ decay mode.
Figure 2: Branching ratios of $\text{Br}(H^{\pm}\rightarrow X)$ in the left panel and $\text{Br}(H^{+}H^{-}\rightarrow XX^{\prime})$ in the right panel. We set $M_{H}=M_{A}=200{\;{\rm GeV}}$.
@@
Let us delve into the decay of the light charged Higgs boson in the mass range between 130 GeV and 170 GeV. There are two dominant decay channels, $H^{\pm}\rightarrow t^{*}b$ and $H^{\pm}\rightarrow\tau\nu$. Figure 2 depicts their branching ratios as functions of $M_{H^{\pm}}$, comparing scenarios with a single charged Higgs boson (left panel) and a pair of charged Higgs bosons (right panel). Note that these results are independent of $\tan\beta$, $m_{12}^{2}$, or $M_{H/A}$. For the $H^{\pm}\rightarrow t^{*}b$ decay, we incorporate QCD radiative corrections to order $\alpha_{s}^{2}$ in the $\overline{\text{MS}}$ scheme, employing the 2HDMC Eriksson:2009ws . This includes adjustments for running fermion masses in the Higgs couplings, applying leading logarithmic corrections across all orders with the renormalization scale $\mu_{R}=M_{H^{\pm}}$.
@@
The decay mode $H^{\pm}\rightarrow t^{*}b$ exhibits significant branching ratios throughout the target $M_{H^{\pm}}$ range and becomes dominant if $M_{H^{\pm}}\gtrsim 135$ GeV. For $M_{H^{\pm}}\lesssim 135$ GeV, $H^{\pm}\rightarrow t^{*}b$ becomes the subleading decay channel, with $H^{\pm}\rightarrow\tau\nu$ being the leading one. In the pair production of charged Higgs bosons, $H^{\pm}\rightarrow t^{*}b$ plays a more significant role. The $H^{+}H^{-}\rightarrow t^{*}b\,t^{*}b$ mode is leading for $M_{H^{\pm}}\gtrsim 143$ GeV. Interestingly, the process $H^{+}H^{-}\rightarrow t^{*}b\tau\nu$ emerges as the most prominent for $M_{H^{\pm}}\lesssim 143$ GeV and remains the second most dominant for $M_{H^{\pm}}\gtrsim 143$ GeV, which benefits from a combinatorial factor of two. In contrast, the $\tau\nu\tau\nu$ final state, heavily emphasized in prior studies for lighter $H^{\pm}$, exhibits markedly reduced branching ratios.
@@
$H^{+}H^{-}$ $\rightarrow t^{*}b\tau\nu.$ (9)
III $H^{\pm}\rightarrow t^{*}b$ at the HL-LHC and a 100 TeV $pp$ collider
@@
where $\tau=\tau^{+},\tau^{-}$, $j$ denotes a light quark jet, and the particles in a square bracket represent decay products originating from the same parent particle. The resultant final state includes two jets, two $b$ jets, a tau lepton, and missing transverse energy. The primary background originates from top quark pair production:
$pp\rightarrow t\bar{t}\rightarrow[bW^{+}][\bar{b}W^{-}]\rightarrow[bjj][b\tau\nu].$ (11)
For the simulation of the signal and background, we followed a comprehensive procedure. We first calculated parton-level cross sections at the 14 TeV LHC using MadGraph5-aMC@NLO Alwall:2011uj version 3.5.0 with the PDF set to NNPDF31_nlo_as_0118, setting both the renormalization and factorization scales as $\mu_{R}=\mu_{F}=\sum_{i}\sqrt{p_{T,i}^{2}+m_{i}^{2}}$. We generated $3.2\times 10^{6}$ signal events and $1.2\times 10^{7}$ background events. Next, we applied NLO corrections for the signal and approximate N3LO corrections for the background by incorporating the $K$-factor Kim:2024ppt . The parton-level cross sections were determined to be $1.593\times 10{\;{\rm fb}}$ for the signal and $1.028\times 10^{2}{\;{\rm pb}}$ for the background.
Parton showering and hadronization were performed using Pythia version 8.309 Bierlich:2022pfr . For the detector-level analysis, a fast detector response simulation was employed with Delphes deFavereau:2013fsa using the high-luminosity LHC card, delphes_card_HLLHC.tcl. Jet clustering was conducted with FastJet version 3.3.4 Cacciari:2011ma using the anti-$k_{T}$ algorithm with a jet radius of $R=0.4$.
@@
  • $E^{\rm miss}_{T}>25$ GeV.
The presence of neutrinos in the decay chain of $H^{\pm}\rightarrow\tau\nu$ necessitates a minimum threshold for the missing transverse energy ${E_{T}^{\rm miss}}$.
@@
Despite the relatively loose selection criteria, the basic selection results in an exceedingly low acceptance rate for the signal, approximately 1%. In contrast, the background acceptance rate is substantially higher, around 2.7%. With the suppressed signal cross section after the basic selection being only $1.852\times 10^{-1}{\;{\rm fb}}$, even the final projected luminosity of $3{\;{\rm ab}^{-1}}$ at the HL-LHC would result in merely a few hundred signal events. This scarcity of signal events severely limits the ability to devise an effective strategy using kinematic cuts to disentangle the signal from the backgrounds, irrespective of their efficiency.
@@
Figure 3: Normalized distributions of transverse momenta for the leading and subleading $b$-jets at the parton level in the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$, with $M_{H^{\pm}}=150{\;{\rm GeV}}$. The left panel shows the distributions at the 14 TeV LHC, while the right panel presents the results for a 100 TeV $pp$ collider.
@@
A primary factor contributing to the low signal acceptance after basic selection is the low transverse momentum (softness) of the $b$ jets originating from the decay $H^{\pm}\rightarrow t^{*}b$. This softness is due to the small mass difference between $M_{H^{\pm}}$ and $m_{t}$, which is crucial for ensuring a substantial branching ratio for the $H^{\pm}\rightarrow t^{*}b$ decay mode, as depicted in Figure 2. Additionally, the $b$ quark from the decay of the off-shell top quark, $t^{*}\rightarrow Wb$, also exhibits lower transverse momentum compared to its on-shell counterpart. The challenge with these soft $b$ jets is that the majority of the subleading $b$ jets fail to meet the minimum transverse momentum threshold ($p_{T}>25{\;{\rm GeV}}$) required for jet clustering.
@@
To illustrate the softness of the $b$ jets, we present in Figure 3 the parton-level $p_{T}$ distributions of the leading and subleading $b$ jets for the signal process. We order jets by their descending $p_{T}$. The left panel shows the results at the 14 TeV LHC, while the right panel displays the results for a 100 TeV $pp$ collider. It is evident that approximately 70% of the subleading $b$ jets fail to surpass the $p_{T}>25{\;{\rm GeV}}$ threshold. This minimum $p_{T}$ threshold for $b$ jets cannot be relaxed because it plays a pivotal role in jet clustering algorithms, primarily aimed at reducing noise from low-energy particles, suppressing background, and enhancing computational efficiency. Moreover, lowering this $p_{T}$ threshold is counterproductive for the signal-to-background analysis, as the background acceptance would increase more than the signal acceptance.
@@
Even at a higher collision energy of 100 TeV, as shown in the right panel, both the leading and subleading $b$ jets remain as soft as those at the HL-LHC. The $b$ jets do not receive a substantial boost because parton-parton collisions at hadron colliders do not fully utilize the beam energy.
@@
Given the limited number of signal events after the basic selection, we need to devise a strategic approach using kinematic cuts that retain as many signal events as possible while effectively suppressing the backgrounds. With this goal in mind, we examined various kinematic distributions and identified key variables that could discriminate the signal from the background.
@@
$\Delta R(b_{1},b_{2})<0.8,\quad\Delta R(j_{i},b_{j})<1.5\quad\text{ for }i,j=1,2.$ (14)
Other crucial discriminating variables pertain to the reconstruction of the charged Higgs boson mass $M_{H^{\pm}}$. In the signal process $H^{+}H^{-}\rightarrow[bbjj][\tau\nu]$, $M_{H^{\pm}}$ can be measured in two complementary ways: through the invariant mass of the $[bbjj]$ system and the transverse mass derived from the $[\tau\nu]$ system. The transverse mass $M_{T}(X)$ is a useful variable defined for a visible particle $X$ and the missing transverse energy $\vec{E}_{T}^{\rm miss}=-\sum_{i}\vec{p}_{T}^{,i}$, where $i$ covers all observed particles:

### v2

@@ 2 Brief review of the light charged Higgs boson in Type-I 2HDM
For the theoretical requirements, we enforce conditions ensuring vacuum stability Ivanov:2008cxa ; Barroso:2012mj ; Barroso:2013awa , a bounded-from-below Higgs potential Ivanov:2006yq , tree-level unitarity in scalar-scalar scatterings Branco:2011iw ; Arhrib:2000is , and perturbativity of the Higgs quartic couplings Chang:2015goa . Furthermore, we require the parameter points to satisfy the current best-fit values for the Peskin-Takeuchi electroweak oblique parameters Peskin:1991sw in the 2HDM framework He:2001tp ; Grimus:2008nb : $S=-0.04\pm 0.10$, $T=0.01\pm 0.12$, and $U=-0.01\pm 0.09$ ParticleDataGroup:2024cfk . The correlations among these oblique parameters are properly incorporated in our analysis. These conditions are evaluated using the public code 2HDMC Eriksson:2009ws . Additionally, we require that the cutoff scale exceed $10{\;{\rm TeV}}$, where the cutoff scale is defined as the energy level at which any of the conditions for tree-level unitarity, perturbativity, or vacuum stability is violated Kim:2023lxc . The evolution of the model parameters via the renormalization group equations is facilitated using the public code 2HDME Oredsson:2018yho ; Oredsson:2018vio .
Regarding experimental constraints, we incorporated measurements at the 95% confidence level, encompassing the Higgs precision data obtained at the LHC, the inclusive $B$-meson decay to $X_{s}\gamma$ Mahmoudi:2009zx ; Arbey:2017gmh ; Sanyal:2019xcp ; Misiak:2017bgg , $B\to K^{*}\gamma$ Belle:2017hum , $B_{s}\to\phi\gamma$ Belle:2014sac , and direct search bounds from LEP, Tevatron, and LHC experiments. For the Higgs precision and direct search constraints, we employed the public code HiggsTools-v1.2 Bahl:2022igd .
@@
In Figure 1, we present the allowed parameter points in the $(M_{H/A},\tan\beta)$ plane for a fixed $M_{H^{\pm}}=150{\;{\rm GeV}}$. The color code represents $m_{12}^{2}$. A key feature of the results is the upper bound on the masses of $H$ and $A$, with $M_{H/A}\lesssim 236{\;{\rm GeV}}$ for $M_{H^{\pm}}=150{\;{\rm GeV}}$. This is a general characteristic of scenarios with a light charged Higgs boson, primarily driven by the stringent theoretical constraints from perturbativity and unitarity. Specifically, for a light $H^{\pm}$, the quartic couplings $\lambda_{3}$ and $\lambda_{4}$ increase significantly with $M_{H/A}$ (see Equation 23). As a result, excessively heavy $M_{H/A}$ values would violate perturbativity and unitarity constraints. Although our main focus in this paper is the light charged Higgs boson, the relatively low upper bounds on $M_{H/A}$ indicate promising discovery prospects for the heavy neutral Higgs bosons at high-energy colliders. Additionally, we observe a lower bound on $\tan\beta$, specifically $\tan\beta\gtrsim 6$.
@@
Now, let us identify the optimal production mechanism at the LHC for the light charged Higgs boson. LHC searches have primarily focused on its production through top quark decay, $t\rightarrow bH^{+}$, followed by $H^{\pm}\rightarrow\tau^{\pm}\nu$ ATLAS:2018gfm ; Sirunyan:2019hkq . Despite comprehensive searches, no new signals have been observed, leading to stringent upper limits on the product of the two branching ratios, $\text{Br}(t\rightarrow bH^{+})\text{Br}(H^{+}\rightarrow\tau^{+}\nu)$. For a given $M_{H^{\pm}}$, $\text{Br}(t\rightarrow bH^{+})$ is determined solely by $\tan\beta$, whereas $\text{Br}(H^{+}\rightarrow\tau^{+}\nu)$ remains constant. As a result, the observed upper bound on their product imposes a strict constraint on $\tan\beta$ for a given $M_{H^{\pm}}$, thereby limiting $\text{Br}(t\rightarrow bH^{+})$. For instance, at $M_{H^{\pm}}=150{\;{\rm GeV}}$ and $\tan\beta=10$, we find that $\text{Br}(t\rightarrow bH^{+})$ is only $7\times 10^{-4}$. Thus, top quark decay is not an effective production channel for $H^{\pm}$ at the LHC.
@@
Given this constraint, alternative production channels for $H^{\pm}$ via the decay of heavier Higgs states, such as $H/A\rightarrow H^{\pm}W^{\mp(*)}$, have been extensively studied Arhrib:2017wmo ; Mondal:2021bxa ; Arhrib:2021xmc ; Biekotter:2023eil . The production of $H$ or $A$ occurs through gluon fusion ($gg\rightarrow H/A$) or associated production ($gg\rightarrow H\rightarrow AZ$, $q\bar{q}\rightarrow Z\rightarrow HZ$). However, these channels present two significant limitations. First, their signal rates strongly depend on model parameters $M_{H/A}$ and $\tan\beta$. Second, as shown in Figure 1, the allowed parameter space forbids on-shell decays for $H/A\rightarrow H^{\pm}W^{\mp}$, preventing us from using the $W$ boson mass constraint to effectively suppress backgrounds.
@@
A more promising production mechanism is the pair production of charged Higgs bosons via the Drell-Yan process. This channel offers a straightforward and model-independent avenue for $H^{\pm}$ production, as the production cross section is exclusively determined by the mass of the charged Higgs boson. Therefore, we focus our analysis on the pair production channel for the $H^{\pm}\rightarrow t^{*}b$ decay mode.
Figure 2: Branching ratios of $\text{Br}(H^{+}\rightarrow X)$ in the upper panel and $\text{Br}(H^{+}H^{-}\rightarrow XX^{\prime})$ in the lower panel. We set $M_{H}=M_{A}=200{\;{\rm GeV}}$. The branching ratios are independent of $\tan\beta$, as all Yukawa couplings of $H^{\pm}$ share a common factor of $1/\tan\beta$, making them dependent only on the fermion masses.
@@
Let us examine the decay patterns of the light charged Higgs boson in the mass range of 130–170 GeV. The dominant decay channels for $H^{+}$ are $t^{*}\bar{b}$ and $\tau^{+}\nu$. Figure 2 shows their branching ratios as functions of $M_{H^{\pm}}$ for two scenarios: a single charged Higgs boson (upper panel) and a pair of charged Higgs bosons (lower panel). These branching ratios are independent of $\tan\beta$, $m_{12}^{2}$, and $M_{H/A}$. Our calculations, performed with 2HDMC Eriksson:2009ws , include QCD radiative corrections up to order $\alpha_{s}^{2}$ in the $\overline{\text{MS}}$ scheme, incorporating running fermion masses in the Higgs couplings and leading logarithmic corrections to all orders at the renormalization scale $\mu_{R}=M_{H^{\pm}}$.
The decay mode $H^{+}\rightarrow t^{*}\bar{b}$ maintains a significant branching ratio across the target $M_{H^{\pm}}$ range and becomes the dominant channel for $M_{H^{\pm}}\gtrsim 135$ GeV. For $M_{H^{\pm}}\lesssim 135$ GeV, however, $H^{+}\rightarrow\tau^{+}\nu$ takes precedence as the leading decay mode, with $H^{+}\rightarrow t^{*}\bar{b}$ remaining subdominant.
@@
In the pair production of charged Higgs bosons, the role of $H^{+}\rightarrow t^{*}\bar{b}$ becomes even more pronounced. The process $H^{+}H^{-}\rightarrow t^{*}\bar{b}\,\bar{t}^{*}b$ dominates for $M_{H^{\pm}}\gtrsim 143$ GeV. Interestingly, the decay channel $H^{+}H^{-}\rightarrow t^{*}b\tau\nu$, which accounts for $H^{+}H^{-}\rightarrow t^{*}\bar{b}\tau^{-}\nu$ and its charge-conjugate process $H^{+}H^{-}\rightarrow\bar{t}^{*}b\tau^{+}\nu$, emerges as the most prominent mode for $M_{H^{\pm}}\lesssim 143$ GeV and remains the second most dominant for $M_{H^{\pm}}\gtrsim 143$ GeV, benefiting from a combinatorial factor of two. In contrast, the $\tau\nu\tau\nu$ final state, which has been the focus of previous studies for the light $H^{\pm}$, exhibits significantly reduced branching ratios.
@@
$H^{+}H^{-}$ $\rightarrow t^{*}b\tau\nu,$ (27)
where $t^{*}b\tau\nu$ accounts for both charge-conjugate processes.
3 $H^{\pm}\rightarrow t^{*}b$ at the HL-LHC and a 100 TeV $pp$ collider
@@
where $\tau=\tau^{+},\tau^{-}$, $j$ denotes a light quark jet, and the particles in a square bracket represent decay products originating from the same parent particle.
The resultant final state consists of two light jets, two $b$ jets, a tau lepton, and missing transverse energy. At the 14 TeV LHC, this final state can arise from several SM processes, including top quark pair production, $Wjjb\bar{b}$, $tbjj$ (single top production in the $t$-channel with an associated jet), $t\bar{t}W$, $t\bar{t}Z$, $t\bar{t}h_{\text{SM}}$, $tWZ$, and $t\bar{t}t\bar{t}$.
To evaluate their relative importance and identify the dominant background contributions, we compute the total cross sections for these processes at leading order using MadGraph5_aMC@NLO Alwall:2011uj version 3.5.0, with the NNPDF31_nlo_as_0118 PDF set:
$\sigma(pp\to t\bar{t})$ $=530.2~{}\text{pb},$ (29)
$\sigma(pp\to W^{\pm}jjb\bar{b})$ $=15.7~{}\text{pb},$
$\sigma(pp\to t\bar{t}b\bar{b}jj)$ $=4.626~{}\text{pb},$
$\sigma(pp\to t\bar{t}W^{\pm})$ $=4.685\times 10^{-1}~{}\text{pb},$
$\sigma(pp\to t\bar{t}Z)$ $=6.103\times 10^{-1}~{}\text{pb},$
$\sigma(pp\to t\bar{t}h_{\text{SM}})$ $=4.168\times 10^{-1}~{}\text{pb},$
$\sigma(pp\to t\bar{t}W^{\pm}Z)$ $=9.991\times 10^{-2}~{}\text{pb},$
$\sigma(pp\to t\bar{t}t\bar{t})$ $=5.135\times 10^{-3}~{}\text{pb},$
where $\tilde{t}=t,\bar{t}$ and $\tilde{b}=b,\bar{b}$. For processes involving jets and/or $b$ jets, we imposed the kinematic cuts $p_{T}^{j,b}>25~{}\text{GeV}$, $|\eta_{j,b}|<2.5$, and $\Delta R_{j_{i},j_{j}}>0.4$ for $j_{i,j}=j,b$. For cross-validation, we additionally used CalcHEP Belyaev:2012qa to calculate the cross sections for several channels such as $pp\to t\bar{t}W^{\pm}$ and $pp\to Wjjb\bar{b}$. The resulting parton-level cross sections agree within expected statistical and systematic uncertainties, confirming the internal consistency of our simulation setup.
As evident from the computed cross sections, top quark pair production is by far the most significant background, with a rate exceeding those of all other processes by more than an order of magnitude. Accordingly, our analysis primarily focuses on this dominant background channel:
$pp\rightarrow t\bar{t}\rightarrow[bW^{+}][\bar{b}W^{-}]\rightarrow[bjj][b\tau\nu],$ (30)
where $[bjj][b\tau\nu]$ implicitly includes all possible charge-conjugate combinations.
To quantitatively assess both the signal and the dominant $t\bar{t}$ background, we carried out a detailed parton-level simulation at the 14 TeV LHC. Cross sections were computed using MadGraph5_aMC@NLO version 3.5.0 with the NNPDF31_nlo_as_0118 PDF set. The renormalization and factorization scales were dynamically set to $\mu_{R}=\mu_{F}=\sum_{i}\sqrt{p_{T,i}^{2}+m_{i}^{2}}$, where the sum runs over all final-state particles. In total, $3.2\times 10^{6}$ signal events and $1.2\times 10^{7}$ background events were generated for further analysis.
To ensure accurate event modeling, we went beyond the leading-order approximation: the signal was simulated at next-to-leading order (NLO), while the $t\bar{t}$ background was corrected to approximate next-to-next-to-next-to-leading order (aN3LO) using $K$-factors based on state-of-the-art theoretical predictions Nason:1987xz ; Beenakker:1990maa ; Czakon:2011xx ; Kidonakis:2019yji ; Kidonakis:2022hfa ; Kim:2024ppt . The resulting parton-level cross sections for the final state in Equation 30 are $1.593\times 10~{}{\;{\rm fb}}$ for the signal and $1.028\times 10^{2}~{}{\;{\rm pb}}$ for the background.
Parton showering and hadronization were performed using Pythia version 8.309 Bierlich:2022pfr . For the detector-level analysis, a fast detector response simulation was employed with Delphes deFavereau:2013fsa using the high-luminosity LHC card delphes_card_HLLHC.tcl. Jet clustering was conducted with FastJet version 3.3.4 Cacciari:2011ma using the anti-$k_{T}$ algorithm with a jet radius of $R=0.4$.
@@
  • ${E_{T}^{\rm miss}}>25$ GeV.
The requirement of a minimum missing transverse energy (${E_{T}^{\rm miss}}$) is essential since our signal includes a neutrino from the decay $H^{\pm}\rightarrow\tau\nu$.
@@
Despite the relatively loose selection criteria, the basic selection results in an exceedingly low acceptance rate for the signal, approximately 1%. In contrast, the background acceptance rate is substantially higher, around 2.7%. With the suppressed signal cross section after the basic selection being only $1.852\times 10^{-1}{\;{\rm fb}}$, even the total luminosity of $3{\;{\rm ab}^{-1}}$ at the HL-LHC would result in merely a few hundred signal events. This scarcity of signal events severely limits the ability to devise an effective strategy using kinematic cuts to disentangle the signal from the backgrounds, irrespective of their efficiency.
@@
Figure 3: Normalized distributions of transverse momenta for the leading and subleading $b$-jets at the parton level in the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$, with $M_{H^{\pm}}=150{\;{\rm GeV}}$. The left panel shows the distributions at the 14 TeV LHC, while the right panel presents the results for a 100 TeV $pp$ collider (solid lines) and a 500 GeV $e^{+}e^{-}$ collider (dashed lines).
@@
A primary factor contributing to the low signal acceptance after basic selection is the low transverse momentum (softness) of the $b$ jets originating from the decay $H^{\pm}\rightarrow t^{*}b$. This softness is due to the small mass difference between $M_{H^{\pm}}$ and $m_{t}$, which is crucial for ensuring a substantial branching ratio for the $H^{\pm}\rightarrow t^{*}b$ decay mode, as depicted in Figure 2. Additionally, the $b$ quark from the decay of the off-shell top quark, $t^{*}\rightarrow Wb$, also exhibits lower transverse momentum compared to its on-shell counterpart. The challenge with these soft $b$ jets is that the majority of the subleading $b$ jets fail to meet the minimum transverse momentum threshold ($p_{T}>25{\;{\rm GeV}}$) required for $b$ jet clustering. [^3] Current $b$-tagging methods at ATLAS and CMS typically require a minimum jet $p_{T}$ threshold of about 20 GeV ATLAS:2019bwq ; CMS:2017wtu , while the FCC-hh studies assume thresholds above 30 GeV Mangano:2022ukr . Although $b$-tagging algorithms are continually improving, lowering this threshold remains particularly challenging. Soft $b$-jets have shorter flight distances, making them difficult to distinguish from light jets. Furthermore, they suffer from lower track reconstruction efficiency, complicating their identification. As demonstrated in Ref. CMS:2017wtu , decreasing the $p_{T}$ threshold further leads to increased background contamination from misidentified light jets, highlighting an intrinsic limitation of $b$-tagging performance at low $p_{T}$.
@@
Figure 3 presents the parton-level $p_{T}$ distributions of the leading and subleading $b$ jets, ordered by descending $p_{T}$, for the signal process. The left panel shows results at the 14 TeV LHC, while the right panel compares results from a 100 TeV $pp$ collider (solid lines) and a 500 GeV $e^{+}e^{-}$ collider (dashed lines). At all three collider configurations, approximately 70% of the subleading $b$ jets fail to exceed the $p_{T}>25{\;{\rm GeV}}$ threshold. This threshold is essential for both $b$-jet tagging and jet clustering algorithms, serving to reduce low-energy noise, suppress backgrounds, and improve computational efficiency. Furthermore, lowering this threshold would be counterproductive, as it would increase background acceptance more than signal acceptance.
@@
Even at the higher collision energy of 100 TeV (solid lines, right panel), both the leading and subleading $b$ jets remain similarly soft as those at the HL-LHC. The $b$ jets do not gain substantial transverse boosts, despite the significant increase in c.m. energy. This occurs because the steep decline of the parton distribution functions at large momentum fractions prevents parton-level collisions at hadron colliders from fully utilizing the available beam energy.
One might wonder whether future $e^{+}e^{-}$ colliders could provide sufficient boost, which is crucially determined by their c.m. energies. Circular colliders such as FCC-ee TLEPDesignStudyWorkingGroup:2013myl ; FCC:2018evy and CEPC An:2018dwb are constrained by synchrotron radiation, limiting them to $\sqrt{s}\leq 365$ GeV Naseem:2019iua . The International Linear Collider (ILC) Moortgat-Pick:2015lbx ; LCCPhysicsWorkingGroup:2019fvj ; Bhattacharya:2023mjr is planned for maximum c.m. energies of 500 GeV, while the Compact Linear Collider (CLIC) Aicheler:2012bya has initial [^4] Although the CLIC aims for a final goal of $\sqrt{s}=3$ TeV, achieving this would require a 50 km tunnel, making it a distant prospect. c.m. energies of 380 GeV.
Since 500 GeV represents the highest achievable c.m. energy for near-future $e^{+}e^{-}$ colliders, we examine the $p_{T}$ distributions of $b$ jets in our signal process $e^{+}e^{-}\to H^{+}H^{-}\\ \to bbjj\tau\nu$ at $\sqrt{s}=500{\;{\rm GeV}}$ (dashed lines in the right panel of Figure 3). These distributions closely resemble those at a 100 TeV $pp$ collider, demonstrating that even at $e^{+}e^{-}$ colliders, the $b$ jets—especially the subleading one—remain insufficiently boosted. While $e^{+}e^{-}$ colliders offer excellent physics opportunities and discovery potential for various new particles, their currently proposed energy range falls short of what is needed to effectively probe our signal process.
@@
Given the small signal yield after basic selection, we focus on developing kinematic cuts that maximize signal retention while effectively reducing backgrounds. With this goal in mind, we examined various kinematic distributions and identified key variables that could discriminate the signal from the background.
@@
$\Delta R(b_{1},b_{2})<0.8,\,\Delta R(j_{i},b_{j})<1.5\,\text{ for}i,j=1,2.$ (33)
Other crucial discriminating variables pertain to the reconstruction of the charged Higgs boson mass $M_{H^{\pm}}$. In the signal process $H^{+}H^{-}\rightarrow[bbjj][\tau\nu]$, $M_{H^{\pm}}$ can be measured in two complementary ways: through the invariant mass of the $[bbjj]$ system and the transverse mass derived from the $[\tau\nu]$ system. The transverse mass $M_{T}(X)$ is a useful variable defined for a visible particle $X$ and the missing transverse energy $\vec{E}_{T}^{\rm miss}=-\sum_{i}\vec{p}_{T}^{,i}$, where $i$ covers all observed particles Smith:1983aa ; Han:2005mu :